OpenStudy (anonymous):

Down at Bikini Bottom Spongebob is 90% water and weighs 10 ounces. On dry land, he is only 80% water, how much does Spongebob weigh now on dry land? Show your solution and why it does or does not make sense.

OpenStudy (anonymous):

if 90%=10 then 10%=10/9 so 80% = 90% -10% = 10 - 1.1 = 8.9 oz As for making sense, hes a talking sponge.....

OpenStudy (anonymous):

okay i get it but how did you get the -1.1?

OpenStudy (radar):

What is missing is the consideration of the non-water material that spongebob is comprised of, and does not change. The 1.1 was the result of dividing 10 by 9.